a gas occupies a volume of 300cm^3 at the pressure of 103,97kpa and at the temperature of 47°c. what volume would the gas occupy at standard condition
Answer #1

Answer-

Initial volume of the gas(V1) = 300 cm3

Initial pressure of the gas(p1) = 103,97 kpa

Initial temperature of the gas(T1) = 47°C = 273.15 + 47 =  320.15 K

At standard condition-

Pressure (P2) = 101.3kPa

Temperatue = 273.15 K

Let volume at the standard condition = V2

Now by ideal gas law-

P1 V1 / T1 = P2 V2 / T2

Put all the know values in the above equation-

103,97 kpa x 300 cm3  / 320.15 K = 101.3kPa x V2 / 273.15 K

\Rightarrow

V2 = (103,97 kpa x 300 cm3  x 273.15 K) / (320.15 K x 101.3kPa)

V2 = 26270.45 cm3